Mr. Meadows Cookie Company (MCC) makes a variety of chocolate chip cookies in a certain plant. Based on orders received and forecasts of buying habits, it is estimated that the demand for the next four months is 850, 1260, 510, and 980, expressed in thousands of cookies. During a 46-day period when there were 120 workers, the company produced 1.7 million cookies. Assume that the number of workdays over the four months is respectively 26, 24, 20, and 16. There are currently 100 workers employed, and there is no starting inventory of cookies.

What is the minimum constant workforce required to meet demand over the next four months with no shortages?

Assume holding cost per unit per month = 10 cents per cookie per month, hiring cost per worker = $100, and firing cost per worker = $200, evaluate the cost of the level strategy plan with a constant workforce as derived in part (a).

Determine the chase demand plan (i.e. the zero inventory plan) that hires and fires workers each month to match demand as closely as possible and the cost of that plan.
Answer #1

Cookies production per person per day = 1.7x1000000 / 120x46 =307.97

[ It is assumed that there was no gap between these workdays]

Demand over next four months = 2750

Max. demand during these months is in month 2, which has 24 working days

Productivity needed per day = 1260000/24 =

Minimum Number of workers needed = 52500/308 = 170.45 =171

This is when there is no inventory is allowed to be kept.

To meet the demand with inventory, the total avg. workforce = Amount of work to be done / productivity per worker

= 2750x1000 / 308 x( 26+24+20+16)

= 2750000 / 308x86 =103.82 =104

Period 1 2 3 4
Demand 850000 1260000 510000 980000
Workers needed 107 170 83 198
Hiring 7 63 0 115
Firing 0 0 87 0
Starting inventory 0 6856 3548 4828
Endng inventory 6856 3548 4828 572

The workers needed are determined by

Demand / no. of days x productivity per person

in period 1 it is

= 850000/26x308 =106.14, which means 107

Goods produced by 107 workers = 107x26x308 = 856856

Excess produced = 856856-850000 = 6856 units which will reflect in inventory. similarly the quantity can be calculated for subsequent periods

Cost of plan

Hiring - 185x100 = 18500

Firing - 87x200 =17400

Inventory = Avg inventory x holding cost per unit

= 15518 x10 cents = 1551.8

Total cost = 18500+17400+1551.8 = 37451.8